# An Improved Ensemble-Based Machine Learning Model with Feature Optimization for Early Diabetes Prediction

### v1

@@ 2 Methodology
This section explains the overall process used in our study, including data collection, preprocessing, feature selection, model development, and evaluation. Each step was carefully designed to ensure accurate, reliable, and interpretable results for diabetes prediction using machine learning.
@@ 2.1 Data Source and Data Preprocessing
The study used the 2015 BRFSS dataset containing 253,680 records with 22 numerical clinical features related to diabetes risk factors (age, glucose levels, BMI, blood pressure, etc.).
Class imbalance was addressed using SMOTE for oversampling minority cases and Tomek Links for undersampling majority cases. Features were normalized to standardize scale, and missing values were handled appropriately. The dataset was split into training (70%) and test (30%) sets for model evaluation.
@@ 2.2 Analysis of Exploratory Data (EDA)
We conducted EDA to understand data distributions, patterns, and quality. Univariate analysis using histograms revealed class imbalance and outliers in continuous and categorical variables. Correlation analysis and VIF identified multicollinearity, with blood pressure, cholesterol, and BMI showing strong associations with diabetes, while other variables proved redundant. These findings guided subsequent feature selection to retain only predictive features.
@@ 2.3 Feature Selection
Three feature selection techniques were used to increase efficiency and decrease dimensionality: LASSO regression, Recursive Feature Elimination (RFE), and Mutual Information (MI). A single table was created by normalising the ranks derived from different techniques. The top 18 predictors were chosen for model training based on the average ranking: [’GenHlth’, ’HighBP’, ’BMI’, ’Age’, ’HighChol’, ’CholCheck’, ’Income’, ’Sex’, ’HeartDiseaseorAttack’, ’HvyAlcoholConsump’, ’AnyHealthcare’, ’DiffWalk’, ’PhysActivity’, ’Smoker’, ’Veggies’, ’Fruits’, ’Education’, ’Stroke’]. This choice lowers the chance of overfitting while striking a balance between explainability and predictive power.
@@ 2.4 Data Splitting
To maintain a consistent class distribution, stratified sampling was employed to split the dataset into 80% training and 20% testing. The test data was set aside for the final assessment to ensure that the findings were objective, while the training data was used to construct models and adjust hyperparameters using k-fold cross-validation.
@@ 2.5 Model Training Strategy
We trained nine supervised learning algorithms ranging from simple models (Logistic Regression, Naive Bayes) to advanced ensembles (Random Forest, XGBoost, LightGBM, CatBoost) to compare explainability against predictive power.
Hyperparameter optimization used a two-step approach: Randomized Search to explore the parameter space broadly, followed by GridSearchCV to fine-tune promising regions for optimal performance. We applied 5-fold cross-validation during training to prevent overfitting and ensure robustness. Final evaluation on the test set used multiple metrics (accuracy, precision, recall, F1-score, ROC-AUC, confusion matrix) for comprehensive model assessment.
@@ 2.6 Development of the Stacking Ensemble Model
We chose to use stacking instead of relying on a single model. This approach combines multiple algorithms to leverage their individual strengths. Looking at Table 1, XGBoost caught our attention. It delivered balanced, strong performance across accuracy, precision, and F1-score. But it had a weakness: recall was relatively low. Why does recall matter? In medical applications, missing even one diabetes case is problematic. We need to identify all positive cases. KNN took a different approach. It excelled at recall—meaning it caught diabetic patients really well. The downside was slightly lower overall accuracy. We realized these two models complemented each other perfectly. So we stacked them together using LightGBM as the meta-learner.
This combination worked well. By blending XGBoost’s balanced performance with KNN’s strong recall, we built a more robust model. It caught more positive cases while maintaining solid accuracy.
@@ 3 Results Analysis
This section presents and discusses the outcomes of our experiments, including the setup, model evaluation, and comparison with previous studies. The goal is to analyze how each model performed, interpret key findings from the results, and highlight the effectiveness of the proposed stacking ensemble model for diabetes prediction.
@@ 3.1 Performance Analysis of All Models
We tested nine machine learning models to find the best approach for diabetes prediction from clinical features. This included traditional methods like Logistic Regression and Naive Bayes, as well as advanced ensemble techniques like Random Forest, XGBoost, LightGBM, and CatBoost.
The stacking ensemble delivered the strongest results, achieving 94.82% accuracy on test data by combining the power of two complementary models. KNN provided exceptional sensitivity (capturing 97.02% of diabetic cases), while XGBoost contributed high precision (93% of positive predictions were correct). By blending these strengths, the stacking approach achieved the best balance across all performance metrics—particularly the highest F1-score of 0.9493, which matters most in medical applications. Naive Bayes was the weakest model with only 71.29% accuracy. Its core weakness is assuming that clinical features are independent when they’re actually interconnected (glucose levels, body weight, and age all influence each other). This fundamental limitation prevented it from capturing the complex patterns needed for accurate diabetes prediction.
@@
Table 1, shows how all nine models performed. The standout finding: ensemble methods significantly outperformed traditional models. Tree-based models (Random Forest, XGBoost, LightGBM, CatBoost) clustered around 90-91% accuracy, while simpler statistical methods lagged below 74%.
@@
Table 2, reveals a critical insight: the stacking model maintains excellent generalization, with only a 3.93% drop from training to test accuracy. In contrast, KNN shows severe overfitting, losing 9.23% performance on unseen data—a key reason why individual KNN couldn’t be relied upon. Ensemble methods consistently maintained stable performance gaps, demonstrating superior generalization compared to traditional statistical approaches.
@@ 3.2 Performance Measure for Ensemble Model
Figure 2, shows the stacking model achieved 36,041 true negatives and 37,589 true positives, with minimal misclassifications (2,821 false positives and 1,273 false negatives), demonstrating balanced diagnostic performance across both classes.
@@
The ROC curve Figure 3, evaluates how well each model distinguishes between diabetic and non-diabetic patients. The stacking ensemble dominates with an AUC of 0.989, meaning it correctly separates the two classes nearly perfectly across all decision thresholds. Random Forest (AUC = 0.968) and LightGBM (AUC = 0.965) also perform strongly, while traditional methods like Naive Bayes (AUC = 0.775) and Logistic Regression (AUC = 0.814) lag significantly behind. This clear separation shows why ensemble methods are superior for clinical diagnostic tasks.
@@
The Precision-Recall curve Figure 7, tells the real-world story for medical applications where catching diabetic cases (recall) while minimizing false alarms (precision) both matter equally. The stacking model achieves an average precision of 0.991, the highest among all models. This exceptional performance means the stacking ensemble reliably identifies diabetic patients while maintaining high confidence in its positive predictions. Random Forest (AP = 0.974) and LightGBM (AP = 0.973) follow closely, but the stacking approach’s 0.991 score confirms its superior reliability for clinical deployment.

### v2

@@ 2 Methodology
This section explains the overall process used in our study, including data collection, preprocessing, feature selection, model development, and evaluation. Each of the steps was designed carefully, ensure reliable and accurate results for diabetes classification using machine learning
@@ 2.1 Data Source and Data Preprocessing
The study used the BRFSS dataset, we assessed a number of supervised learning techniques. Class imbalance was addressed using SMOTE and Tomek Links were used to correct class imbalance. This is used for undersampling the majority of cases. this method normalized the feature and handled missing values correctly. The 2015 BRFSS dataset which includes roughly 253,680 records with 22 numerical features.
@@ 2.2 Analysis of Exploratory Data (EDA)
Prior to modelling, we used EDA to understand data distributions, patterns, and quality. Univariate analysis. Distributions, imbalances, and outliers were observed using histograms for both continuous and categorical variables. Correlation analysis. Correlation analysis and VIF identified multicollinearity, with some variables seemed less significant, others, like blood pressure, cholesterol, and BMI, revealed higher relationships with diabetes.
@@ 2.3 Feature Selection
We used three feature selection techniques to increase efficiency and decrease dimensionality: LASSO regression, Recursive Feature Elimination (RFE), and Mutual Information (MI). A single table was created by normalising the ranks derived from different techniques.
The top 18 predictors were chosen for model training based on the average ranking: [’GenHlth’, ’HighBP’, ’BMI’, ’Age’, ’HighChol’, ’CholCheck’, ’Income’, ’Sex’, ’HeartDiseaseorAttack’, ’HvyAlcoholConsump’, ’AnyHealthcare’, ’DiffWalk’, ’PhysActivity’, ’Smoker’, ’Veggies’, ’Fruits’, ’Education’, ’Stroke’]. This selection lowers the chance of overfitting while striking a balance between explainability and predictive power.
@@ 2.4 Data Splitting
About 80% of the information was used for training and 20% was used for testing. The test data was set aside for the final assessment to ensure that the findings were unbiased, while the training data was used to construct models and adjust hyperparameters using k-fold cross-validation.
@@ 2.5 Model Training Strategy
For our study, we applied several supervised machine learning algorithms for model development. These include logistic regression, random forest, Gradient Boosting, XGBoost, Support Vector Classifier (SVC), LightGBM, CatBoost, Naive Bayes, and K Nearest Neighbors (KNN). The aim was to compare both simple explainable models and advanced ensemble learners to achieve the best outcomes.
For hyperparameter optimisation, we follow a two-step process. In the first step, we apply Randomized Search to broadly explore a wide range of hyperparameter space. GridSearchCV is used to fine-tune the models for optimal performance once the promising parameter has been identified. This approach ensures both computational efficiency and achieves the best performance. To prevent overfitting and ensure robustness, we use 5-fold cross-validation during training. The final model was evaluated on the unseen test sets, where multiple performance metrics were considered, including the classification report, confusion matrix, and ROC-AUC, to provide a fair and balanced assessment of the models.
@@ 2.6 Development of the Stacking Ensemble Model
We chose to use stacking instead of depending on a single model. This approach combines multiple algorithms by capturing their individual strengths. Looking at Table 1, XGBoost caught our attention. It provides balanced and solid performance across accuracy, precision, and F1-score. However, it had a weakness, recall was slightly low, but in medical applications, recall plays a very significant role. Even missing a single diabetes case may cause a problem. That’s why we need to identify all of the positive cases. KNN took a different approach. It emphasizes recall, which means it holds diabetes patients really well. Overall, the weakness is slightly lower accuracy. We realized that two of these models will fulfill each other perfectly, so we stacked them together using LightGBM as a meta learner.
This combination worked well. By mixing XGBoost’s balanced performance with KNN’s strong recall, we built a more robust model. It holds more positive cases with maintains a solid accuracy.
@@ 3 Results Analysis
In this section, we will illustrate and discuss about the outcomes of our experiments. including setup, model evaluation, contribution, and comparison with the previous works. The goal is to analyze how each model performed, explain the key findings from the results. and highlight how our proposed stacking ensemble model overperformed for diabetes prediction.
@@ 3.1 Performance Analysis of All Models
In this study, we used nine different types of machine learning models to predict diabetes based on features. We used supervised machine learning algorithms, including Logistic Regression and Naive Bayes, which performed the lowest. Tree-based algorithms such as Random Forest, XGBoost, LightGBM, and CatBoost performed much better, with around 90-91% accuracy. Among them, XGBoost provided the best precision, and KNN provided the highest recall. Finally, stacking ensemble model with XGBoost and KNN, which achieved the best outcome. This model achieved the highest accuracy of 94.82% and also gave the highest F1-score and ROC-AUC. The stacking method worked well because XGBoost is used for accuracy, and KNN is used for strong recall. Overall, the ensemble model outperformed any single model.
@@
Table 1, demonstrates the performance of all nine models, where tree-based models performed much better. In addition, ensemble methods significantly outperformed traditional models.
@@
Table 2, demonstrated a significant point. It showed in the stacking model that the difference between training and testing accuracy is low, which is why it performed well on new data. But KNN provided the best performance in training, on the other hand, in testing, it performed poorly. So KNN showed overfitting individually. By contrast, the ensemble model did not show this problem and was more stable compared to traditional statistical approaches.
@@ 3.2 Performance Measure for Ensemble Model
Figure 2, shows the stacking model achieved 36,041 true negatives and 37,589 true positives, with few misclassifications: 2,821 false positives and 1,273 -neg), while performing with balance on both. classes.
@@
ROC curve Figure 3, shows the capability of each model in differentiating between diabetic and non-diabetic patients. The stacking ensemble dominates with an AUC of 0.989 - i.e., it separates the two classes nearly perfectly. across all decision thresholds. Random Forest (AUC = 0.968) and LightGBM AUC = 0.965), while traditional methods such as Naive Bayes also perform well. While LDA, SVM, and Logistic Regression are far behind at AUCs of 0.775 and 0.814, respectively.
@@
The clear separation reflects why ensemble methods are superior for the clinical task at hand. Precision-Recall Curve Figure 4 tells the real-world story. for medical applications which involve catching diabetic cases, recall while minimizing Both false alarms and precision are important. The stacking model realizes an an average precision of 0.991, which is the best for all models. This excellent performance indicates that the stacking ensemble reliably identifies diabetic patients while high confidence for the positive predictions it makes. Random Forest (AP = 0.974) and LightGBM (AP = 0.973) follow closely, but the stacking approach’s 0.991 score confirms its superior reliability for clinical deployment.